# Towards Explainable Fake Image Detection with Multi-Modal Large Language Models

### v1

@@ Abstract.
Recent advancements in image generation have provoked social and security concerns, yet most detection methods rely on black-box models that generalize poorly. By utilizing advances in Multi-modal Large Language Models (MLLMs), we propose a framework that fuses six specialized paradigms, each analyzing a distinct aspect of the image, to provide a final verdict with coherent, evidence-based reasoning. Experimental results on a diverse dataset of real and AI-generated images demonstrate that our approach outperforms both traditional detection methods and top humans, while providing . This study underscores the potential of MLLMs in developing robust, explainable, and reasoning-driven detection systems. The code is available at https://github.com/Gennadiyev/mllm-defake.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; Oct 27, 2025; Dublin, Irelandisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Figure 1. Traditional methods do not provide reasons for their predictions, while our method provides sensible reasons behind the verdict.
[…]
An image of a plate of spaghetti, when provided as input to traditional methods, e.g. CNNSpot, we can obtain a verdict that the image is AI-generated. Our MLLM-powered method is able to provide natural language rationale behind the decision.
@@
Detecting AI-generated images remains a challenging task, usually formulated as a binary classification problem. While deep learning-based methods have exhibited solid performance (Chen et al., 2024a; Tan and Le, 2019), they are prone to overfitting and often rely heavily on large, labeled datasets of real and synthetic images. Alternative approaches focus on statistical anomalies (Ricker et al., 2024; Bi et al., 2023) but frequently struggle with poor generalization, especially when limited to specific generative models or image domains (e.g., facial images). Consequently, there is a pressing need for detection models that not only perform well across diverse generative techniques but also provide interpretability, enabling users to understand why images are flagged.
With the progress of multi-modal large language models (MLLMs), researchers have begun exploring their potential in detecting AI-generated images. A key advantage of using MLLMs is their ability to generate predictions with explanations, thus enhancing interpretability. Studies have shown that an open-source MLLM (mPlug-owl2 (Ye et al., 2024)) with 7B parameters can achieve an accuracy of 71.78%, while OpenAI’s GPT-4V reaches 78.03%, outperforming the average human accuracy of 74.51% (Li et al., 2024). Zhang et al. (2024c) highlights the quality of reasons that MLLMs provide when verifying image authenticity. Although these models demonstrate notable potential, MLLMs remain susceptible to hallucination (Gao et al., 2024) and generally cannot surpass the accuracy of traditional methods. Meanwhile, their responses can be sensitive to prompt design (Wu et al., 2024), suggesting the need for more refined techniques to enhance both accuracy and reasoning quality.
In this work, we enhance the capabilities of multi-modal large language models (MLLMs) through an innovative interrogation methodology. By integrating human expertise with MLLM-based reasoning, we design six distinct prompts, each focusing on a distinct aspect of the image. By fusing the model responses from these prompts, we achieve a higher overall accuracy and improved interpretability. Our quantitative and qualitative analysis reveals that MLLMs can process visual information similarly to human perception when distinguishing between real and AI-generated images. The models exhibit strong generalization capabilities, with GPT-4o achieving 93.4% accuracy, outperforming traditional deep learning models (Chen et al., 2024a) (91.8%) and AEROBLADE (Ricker et al., 2024) (85.2%), as well as the most accurate human annotator (86.3%).
@@
Figure 2. The overall design of the proposed MLLM-based AI-generated image detection framework.
[…]
Sample output from all six prompts and fusion stage. Text is based on two images, an AI-generated dog image, and a natural photograph of a ring.
2. Related Work
@@ 2.1. AI-Generated Image Detection Methods
Beyond artifact-based methods, deep-learning based methods are also advancing. ResNet and Vision Transformers trained on real and synthetic samples can effectively distinguish real from AI-generated images (Chen et al., 2024a; Tan and Le, 2019; Park and Owens, 2024). This paradigm leverages deep networks’ strong feature extraction capabilities to automatically learn discriminative patterns. However, the effectiveness of large-scale training heavily depends on data diversity and quality. Generalization to unseen generative models, particularly from GANs to diffusion-based methods, remains challenging. Meanwhile, as diffusion models produce increasingly photorealistic images, traditional artifact-based cues, including unnatural hues, faulty perspectives, and blurry text, are no longer reliable. Additionally, Grommelt et al. (2024) highlighted that existing models are prone to overfitting on JPEG artifacts, further limiting their robustness.
@@ 2.2. Multi-Modal Large Language Models
Multi-modal large language models (MLLMs) are large language models (LLMs) enhanced with vision or audio modalities (Lin et al., 2024; Bavishi et al., 2023). Over the past few years, both open-source and proprietary models have significantly advanced vision-language capabilities. Open-source ones include Qwen-VL (Bai et al., 2023; Wang et al., 2024; Qwen Team, 2025) and QVQ (Qwen Team, 2024), InternVL (Chen et al., 2024b), Ovis (Lu et al., 2024b), LLaVA and its predecessors (Liu et al., 2023b, a, 2024), Llama-3 (The Llama Team, 2024), among many others. Proprietary models have also progressed rapidly. Notable releases include GPT-4o, GPT-4o-mini (OpenAI, 2024b, a) and Gemini-1.5 (Gemini Team, 2024b, a). These models demonstrate remarkable performance in vision-related tasks, including image reasoning and interpretation. Most benchmarks for vision language models focus on reasoning capabilities related to the image (Ying et al., 2024; Liao et al., 2024) or geometric problem-solving skills (Kamoi et al., 2024; Lu et al., 2024a).
Li et al. (2024) proposed FakeBench, a benchmark for general AI-generated image detection. It includes three primary datasets for detection, reasoning, interpretation, and open-question answering. FakeBench covers ten generative models across diverse image content, revealing that current MLLMs demonstrate moderate detection accuracy, basic reasoning abilities, and limited fine-grained authenticity analysis. Zhang et al. (2024c) created A-Bench, which “challenges both high-level semantic understanding and low-level visual quality perception” of MLLMs in the form of multiple-choice questions. The question-answer pairs are annotated by humans. While the questions attempt to point out the major flaws in each image, they cannot inspire the MLLMs to actively generate them.
@@ 3.4. Content-Based Prompts
[What is the main subject in the image? Answer concisely with a short descriptive noun phrase, e.g., ’a red bus’, ’a daisy’.]
@@ 3.5. Fusion Process
Prompts categorized as General, Few-Shot, and Content-Based (P1–P6) enhance AI-generated image detection via MLLMs by targeting interpretability aspects like local inspections (P2, P4), logical consistency (P3, P5), and real-world traits (P1, P6). We propose two fusion methods to determine the final verdict. The primary approach sequentially executes P1–P6 in a single pipeline, consolidating outputs for a prediction that balances accuracy and explainability. Alternatively, a majority vote method tallies individual prompt verdicts, trading some explainability for computational efficiency while remaining effective in most cases. Experimental results validate these trade-offs, showing the sequential method excels in interpretability and the majority vote offers practical efficiency.
@@ 4. Experiments
4.1. Dataset Distribution
Figure 3. Distribution of AI generation methods among images in the dataset.
[…]
A chart showing the distribution of AI generation methods among images in the dataset. The chart is divided into sections representing different methods, including GANs, diffusion models, and others. Each section is labeled with the percentage of images generated by that method. A majority of the images are generated by diffusion models, followed by GANs and other methods.
@@ 4.2. Experimental Setup
We ensure relevance and effectiveness in real-world AI-generated image detection by covering a diverse range of generation models, including but not limited to GANs, auto-regressive models, and diffusion models. To achieve this, we selected 1,000 real images and 1,000 AI-generated images, spanning multiple generation architectures, with detailed distribution provided in the appendix. To benchmark MLLM performance against humans, we invited 24 volunteers to classify the images, 12 of whom had prior experience with AI image generation.
We tested recent MLLMs such as GPT-4o, GPT-4o-mini, Llama 3.2 VI (Llama-3.2-Vision-Instruct 11B) (The Llama Team, 2024), LLaVA-CoT (Xu et al., 2024), Qwen-VL (Bai et al., 2023; Wang et al., 2024; Qwen Team, 2025, 2024), Ovis (Lu et al., 2024b) and InternVL series (Chen et al., 2024b). Among them, we selected four models for our full analyses: GPT-4o, GPT-4o-mini, Llama 3.2 VI, and LLaVA-CoT. In our experiments, we used OpenAI GPT-4o that points to gpt-4o- 2024-08-06 and GPT-4o-mini that points to gpt-4o-mini- 2024-07-18. The Llama 3.2 VI refers to the Llama-3.2-Vision-Instruct 11B model (The Llama Team, 2024). LLaVA-CoT is a chain-of-thought model fine-tuned from Llama-3.2-Vision (Xu et al., 2024). Local MLLMs are deployed using vllm-0.7.2 (Kwon et al., 2023) on four NVIDIA A100-40G GPUs.
To compare the performance with traditional methods, we select AEROBLADE (Ricker et al., 2024), CNNSpot (Chen et al., 2024a), CommunityForensics (Park and Owens, 2024) and ObjectFormer (Wang et al., 2022). We evaluate them on the same sampled dataset used for MLLM and human evaluations. For AEROBLADE, we employ Stable Diffusion versions 1.5, 2.1, and 3.5-large (Rombach et al., 2022; Esser et al., 2024) as reconstruction models and select SD3.5-Large for the results, since it provides the best performance. For CNNSpot and ObjectFormer, since the pretrained models performed suboptimally, we evaluated on checkpoints that we trained on the same data. CommunityForensics requires no training.
Figure 4. (a) Examples where GPT-4o fusion gives correct result, while CNNSpot, AEROBLADE, and vanilla GPT-4o (P0) fail. (b) A fusion example. GPT-4o can combine responses from P1-6 effectively, drawing conclusions from reasons instead of verdicts.
[…]
A detailed figure showing two parts: (a) examples of GPT-4o fusion correctly identifying cases where CNNSpot, AEROBLADE, and vanilla GPT-4o fail; and (b) excerpt of text generated from the fusion process where GPT-4o combines responses from multiple prompts P1-6, focusing on reasoning rather than just verdicts.
@@ 4.3. AI-Generated Image Detection Accuracy
The results are summarized in Table 1, reporting accuracy for MLLMs: GPT-4o, GPT-4o-mini, Llama 3.2 VI, LLaVa-CoT; traditional methods: AEROBLADE (Ricker et al., 2024), CNNSpot (Chen et al., 2024a), CommunityForensics (Park and Owens, 2024), DMimageDetection (Corvi et al., 2023) and NPR (Ojha et al., 2023); and human performance. For MLLMs, we report results for baseline (P0), all proposed prompts, majority vote of P1-P6 (Maj.), and fused results. Humans performed well overall (81.9% accuracy), particularly on real images (84.8% accuracy). The best volunteer achieved 4.4% higher accuracy than the average human performance and slightly outperformed the best MLLM baseline (86.3% vs. 85.2% for GPT-4o P0). Other MLLMs performed worse than GPT-4o. While CNNSpot (91.8%) surpassed GPT-4o (85.2%), it operates as a black-box model.
@@ 4.4. Prompt Effectiveness
Figure 5. Percentage of cases where prompts give different verdicts. The data is aggregated from all four models evaluated.
[…]
A six by six grid showcasing the conflict rates of six prompts in a grid.
@@ 4.6. Discussions on MLLM Rejections
MLLMs may generate non-parsable responses or refuse to comment on the given image, termed “rejections,” when processing certain images. The rejection rate varies across different models, and the types of images rejected differ accordingly. A frequent justification for these rejections by MLLMs is their inability to provide commentary on images, particularly those depicting faces, due to ethical considerations. These rejections can be mitigated by modifying the prompt, leading to lower rejection rates. As shown in Table 3, for the 2,000 images evaluated, replacing the word “fake” with “generated” resulted in a 0.95% (19 samples) reduction in rejections for GPT-4o and 1.3% for GPT-4o-mini. In comparison, open-source models generally exhibit lower rejection rates. Open-source models such as Llama-3.2-Vision-Instruct and LLaVA-CoT have a much lower rejection rate than GPT-4o and -4o-mini.

### v2

@@ Abstract.
Recent advancements in image generation have provoked social and security concerns, yet most detection methods rely on black-box models that generalize poorly. By utilizing advances in Multi-modal Large Language Models (MLLMs), we propose a framework that fuses six specialized paradigms, each analyzing a distinct aspect of the image, to provide a final verdict with coherent, evidence-based reasoning. Experimental results on a diverse dataset of real and AI-generated images demonstrate that our approach outperforms both traditional detection methods and top humans, while providing explainability. This study underscores the potential of MLLMs in developing robust, explainable, and reasoning-driven detection systems. The code is available at https://github.com/Gennadiyev/mllm-defake.
Generative AI, Multi-modal Large Language Models, Explainable AI, AI safety
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755421isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Natural language generationccs: Computing methodologies Computer visionccs: Social and professional topics Censorship
@@ 1. Introduction
Figure 1. Traditional methods do not provide reasons for their predictions, while our method provides sensible reasons behind the verdict.:
Figure 2. The overall design of the proposed MLLM-based AI-generated image detection framework.:
@@
Detecting AI-generated images remains a challenging task, usually formulated as a binary classification problem. While deep learning-based methods have exhibited solid performance (Wang et al., 2020; Tan and Le, 2019), they are prone to overfitting and often rely heavily on large, labeled datasets of real and synthetic images. Alternative approaches focus on statistical anomalies (Ricker et al., 2024; Bi et al., 2023) but frequently struggle with poor generalization, especially when limited to specific generative models or image domains (e.g., facial images). Consequently, there is a pressing need for detection models that not only perform well across diverse generative techniques but also provide interpretability, enabling users to understand why images are flagged.
With the development of multi-modal large language models (MLLMs), researchers have begun exploring their potential in detecting AI-generated images. A key advantage of using MLLMs is their ability to generate predictions with explanations, thus enhancing interpretability. Studies have shown that an open-source MLLM (mPlug-owl2 (Ye et al., 2024)) with 7B parameters can achieve an accuracy of 71.78%, while OpenAI’s GPT-4V reaches 78.03%, outperforming the average human accuracy of 74.51% (Li et al., 2024). Zhang et al. (2024c) highlights the quality of reasons that MLLMs provide when verifying image authenticity. Although these models demonstrate notable potential, MLLMs remain susceptible to hallucination (Gao et al., 2024) and generally cannot surpass the accuracy of traditional methods. Meanwhile, their responses can be sensitive to prompt design (Wu et al., 2024), suggesting the need for more refined techniques to enhance both accuracy and reasoning quality.
In this work, we enhance the capabilities of multi-modal large language models (MLLMs) through an innovative interrogation methodology. By integrating human expertise with MLLM-based reasoning, we design six distinct prompts, each focusing on a distinct aspect of the image. By fusing the model responses from these prompts, we achieve a higher overall accuracy and improved interpretability. Our quantitative and qualitative analysis reveals that MLLMs can process visual information similarly to human perception when distinguishing between real and AI-generated images. The models exhibit strong generalization capabilities, with GPT-4o achieving 93.4% accuracy, outperforming traditional deep learning models (Wang et al., 2020) (91.8%) and AEROBLADE (Ricker et al., 2024) (85.2%), as well as the most accurate human annotator (86.3%).
@@
2. Related Works
@@ 2.1. AI-Generated Image Detection Methods
Beyond artifact-based methods, deep-learning based methods are also advancing. ResNet and Vision Transformers trained on real and synthetic samples can effectively distinguish real from AI-generated images (Wang et al., 2020; Tan and Le, 2019; Park and Owens, 2024). This paradigm leverages deep networks’ strong feature extraction capabilities to automatically learn discriminative patterns. However, the effectiveness of large-scale training heavily depends on data diversity and quality. Generalization to unseen generative models, particularly from GANs to diffusion-based methods, remains challenging. Meanwhile, as diffusion models produce increasingly photorealistic images, traditional artifact-based cues, including unnatural hues, faulty perspectives, and blurry text, are no longer reliable. Additionally, Grommelt et al. (2024) highlighted that existing models are prone to overfitting on JPEG artifacts, further limiting their robustness.
@@ 2.2. Multi-Modal Large Language Models
Multi-modal large language models (MLLMs) are large language models (LLMs) enhanced with vision or audio modalities (Lin et al., 2024; Bavishi et al., 2023). Over the past few years, both open-source and proprietary models have significantly advanced vision-language capabilities. Open-source ones include Qwen-VL (Bai et al., 2023; Wang et al., 2024; Qwen Team, 2025) and QVQ (Qwen Team, 2024), InternVL (Chen et al., 2024), Ovis (Lu et al., 2024b), LLaVA and its predecessors (Liu et al., 2023b, a, 2024), Llama-3 (The Llama Team, 2024), among many others. Proprietary models have also progressed rapidly. Notable releases include GPT-4o, GPT-4o-mini (OpenAI, 2024b, a) and Gemini-1.5 (Gemini Team, 2024b, a). These models demonstrate remarkable performance in vision-related tasks, including image reasoning and interpretation. Most benchmarks for vision-language models focus on reasoning capabilities related to the image (Ying et al., 2024; Liao et al., 2024) or geometric problem-solving skills (Kamoi et al., 2024; Lu et al., 2024a).
Li et al. (2024) proposed FakeBench, a benchmark for general AI-generated image detection that includes three primary datasets: reasoning, interpretation, and open-question answering. It covers ten generative models across diverse image content, revealing that current MLLMs demonstrate moderate detection accuracy, basic reasoning abilities, and limited fine-grained authenticity analysis. Zhang et al. (2024c) created A-Bench, which “challenges both high-level semantic understanding and low-level visual quality perception” of MLLMs in the form of multiple-choice questions, pointing out major flaws in each AI-generated image. The question-answer pairs are curated by humans and not actively generated by MLLMs. FakeReasoning (Gao et al., 2025) conducts the forgery reasoning task in structured steps using MLLMs, giving promising results by fine-tuning LLaVA-1.5-13b on the MMFR-Dataset, created using GPT-4o and human experts.
@@ 3.4. Content-Based Prompts
[What is the main subject in the image? Please answer concisely with a short descriptive noun phrase, e.g., ’a red bus’, ’a daisy’.]
@@ 3.5. Fusion Process
Prompts categorized as General, Few-Shot, and Content-Based enhance AI-generated image detection via MLLMs by targeting interpretability aspects like local inspections (P2, P4), logical consistency (P3, P5), and real-world traits (P1, P6). We propose two fusion methods to determine the final verdict. The primary approach executes P1–P6 in parallel before consolidating an output, balancing accuracy and explainability. Alternatively, a majority vote method tallies individual prompt verdicts, trading some explainability for computational efficiency while remaining effective in most cases. Experimental results validate these trade-offs, showing the sequential method excels in interpretability and the majority vote offers practical efficiency.
@@ 4.1. Experimental Setup
We ensure relevance and effectiveness in real-world AI-generated image detection by covering a diverse range of generation models, including but not limited to GANs, auto-regressive models, and diffusion models. To achieve this, we selected 1,000 real images and 1,000 AI-generated images from WildFake (Hong et al., 2025), spanning different generation architectures, with detailed distribution provided in the appendix. To benchmark MLLM performance against humans, we invited 24 volunteers to classify the images, 12 of whom had prior experience with AI image generation.
We tested recent MLLMs such as GPT-4o, GPT-4o-mini, Llama 3.2 VI (Llama-3.2-Vision-Instruct 11B) (The Llama Team, 2024), LLaVA-CoT (Xu et al., 2024), Qwen-VL (Bai et al., 2023; Wang et al., 2024; Qwen Team, 2025, 2024), Ovis (Lu et al., 2024b) and InternVL series (Chen et al., 2024). Among them, we selected four models for our full analyses: GPT-4o, GPT-4o-mini, Llama 3.2 VI, and LLaVA-CoT. In our experiments, we used OpenAI GPT-4o that points to gpt-4o- 2024-08-06 and GPT-4o-mini that points to gpt-4o-mini- 2024-07-18. The Llama 3.2 VI refers to the Llama-3.2-Vision-Instruct 11B model, from which LLaVA-CoT is a fine-tuned chain-of-thought model. Local MLLMs are deployed using vllm-0.7.2 (Kwon et al., 2023) on four NVIDIA A100-40G GPUs.
To compare the performance with traditional methods, we select AEROBLADE (Ricker et al., 2024), CNNSpot (Wang et al., 2020), CommunityForensics (Park and Owens, 2024) and ObjectFormer (Wang et al., 2022). We evaluate them on the same sampled dataset used for MLLM and human evaluations. For AEROBLADE, we employ Stable Diffusion versions 1.5, 2.1, and 3.5-large (Rombach et al., 2022; Esser et al., 2024) as reconstruction models and select SD3.5-Large for the results, since it provides the best performance. For CNNSpot and ObjectFormer, since the pretrained models performed suboptimally, we evaluated on checkpoints that we trained on the same data. CommunityForensics requires no training.
Figure 3. (a) Examples where GPT-4o fusion gives correct result, while CNNSpot, AEROBLADE, and vanilla GPT-4o (P0) fail. (b) A fusion example. GPT-4o can combine responses from P1-6 effectively, drawing conclusions from reasons instead of verdicts.:
@@ 4.2. AI-Generated Image Detection Accuracy
The results are summarized in Table 1, reporting accuracy for MLLMs: GPT-4o, GPT-4o-mini, Llama 3.2 VI, LLaVa-CoT; traditional methods: AEROBLADE (Ricker et al., 2024), CNNSpot (Wang et al., 2020), CommunityForensics (Park and Owens, 2024), DMimageDetection (Corvi et al., 2023) and NPR (Ojha et al., 2023); and human performance. For MLLMs, we report results for baseline (P0), all proposed prompts, majority vote of P1-P6 (Maj.), and fused results. Humans performed well overall (81.9% accuracy), particularly on real images (84.8% accuracy). The best volunteer achieved 4.4% higher accuracy than the average human performance and slightly outperformed the best MLLM baseline (86.3% vs. 85.2% for GPT-4o P0). Other MLLMs performed worse than GPT-4o. While CNNSpot (91.8%) surpassed GPT-4o (85.2%), it operates as a black-box model.
@@ 4.3. Prompt Effectiveness
Figure 4. Percentage of cases where prompts give different verdicts. The data is aggregated from all four models evaluated.:
@@ 4.5. Discussions on MLLM Rejections
MLLMs may generate non-parsable responses or refuse to comment on the given image, termed “rejections,” when processing certain images. The rejection rate varies across different models, and the types of images rejected differ accordingly. A frequent justification for these rejections by MLLMs is their inability to provide commentary on images, particularly those depicting faces, due to ethical considerations. These rejections can be mitigated by modifying the prompt, leading to lower rejection rates. Table 3 shows that among the 2,000 images evaluated, replacing the word “fake” with “generated” resulted in a 0.95% (19 samples) reduction in rejections for GPT-4o and 1.3% for GPT-4o-mini. In comparison, open-source models generally exhibit lower rejection rates. Open-source models such as Llama-3.2-Vision-Instruct and LLaVA-CoT have a much lower rejection rate than GPT-4o and -4o-mini.
@@ 6. Conclusion
Acknowledgements.
This research is supported, in part, by the National Natural Science Foundation of China (Grant No. 62302295), by Ant Group, the Shanghai Municipal Science and Technology Major Project, China (Grant No. 2021SHZDZX0102), the Pioneer R&D Program of Zhejiang Province (No. 2024C01024) and the foundation of Key Laboratory of Artificial Intelligence, Ministry of Education, P.R. China.
Appendix
Prompt GPT-4o GPT-4o-mini Gemini 1.5 Pro Kimi-VL-A3B-Instruct
P0 0.852 0.855 0.849 0.769 0.910 0.628 0.758 0.906 0.610 0.585 0.622 0.548
MiMo-VL-7B Qwen2-VL-7B-Instruct QVQ-72B-Preview Qwen2.5-VL-7B-Instruct
P0 0.585 0.887 0.283 0.623 0.901 0.346 0.643 0.789 0.497 0.621 0.660 0.582
Ovis1.6-Gemma2-9B InternVL2-8B-MPO InternVL2.5-26B MiniCPM-V-2.6 (8B)
P0 0.559 0.828 0.290 0.592 0.623 0.561 0.566 0.673 0.460 0.545 0.766 0.324
Table 4. Experimental result of P0 on different open-source and proprietary models.
Appendix A Dataset Distribution
Figure 5. Distribution of AI generation methods among images in the dataset.:
Appendix B Discussions on Model Responses
To accommodate paper length constraints, we present additional results for qualitative analysis in Figure 9. These examples are unfiltered, showcasing the model’s typical performance. Our analysis shows that GPT-4o produces varying verdicts across different prompts but maintains consistent reasoning accuracy during the fusion stage. This aligns with our experimental findings, where post-fusion responses achieve the highest accuracy among MLLM-based approaches. Our full prompts can be accessed from this link.
For conciseness, responses are further shortened with the prompt below.
<original_response>
Summarize the above response into no more than 5 key points, where each point should be fewer than 6 words. Output a numbered list in markdown format. Do not include the final verdict in your response.
Figure 6. Common key words present in model responses.:
Key words that model uses. To qualitatively assess the model’s attention tendency, we list the most common key words that are seen many times in model responses. While some of the model responses
Appendix C Experimental Results on Other MLLMs
We evaluate several MLLMs to distinguish AI-generated images from real ones. The results, as shown in Table 4, indicate that Qwen2-VL (Wang et al., 2024) and QVQ (Qwen Team, 2024) perform relatively well in this task. However, Qwen2-VL exhibits a bias towards classifying images as real, while QVQ, with 72B parameters, can take almost ten minutes for one fusion query on 4x NVIDIA H800 GPU. Qwen2.5-VL-7B-Instruct also gives remarkable results, but it lacks the multi-turn image input chat template, rendering P2 and P4 unusable. Gemini 1.5 Pro (Gemini Team, 2024a), as a proprietary model, scored lower than GPT-4o-mini in all three metrics. We selected LLaVA-CoT (Xu et al., 2024) for its reinforced CoT capabilities and conducted further experiments to explore the potential of open-source MLLMs in AI-generated image detection. Llama 3.2 Vision Instruct, the base model of LLaVA-CoT, is used for comparison.
Figure 7. Grid representation of accuracy changes for different real-fake exemplar pairs in the few-shot learning setting. Positive values indicate accuracy improvements over the zero-shot baseline (P-Zero), while negative values denote declines.:
Figure 8. Impact of real and fake sample selection on the overall accuracy of GPT-4o. The dashed horizontal line score represents zero-shot accuracy, while the solid horizontal line line denotes the average accuracy across all two-shot pairs.:
Appendix D Discussions on Few-Shot Exemplar Selection
MLLMs exhibit sensitivity to exemplar selection in few-shot learning (Guo et al., 2024). To systematically assess its impact, we compare few-shot learning (P3) against the zero-shot paradigm (P0), evaluating how exemplar choice affects model accuracy and decision stability.
We selected six real and six AI-generated images uniformly. These images were independently annotated by human researchers, identifying artifacts, inconsistencies, and perceptual anomalies that might influence MLLM classification. All 36 possible real-fake combinations were tested on the same dataset as used in Section 4. Additionally, we removed both exemplars from the prompt to create a zero-shot context (P-Zero), serving as the control group.
Figure 8 presents the comparative accuracy of different few-shot sample combinations against the zero-shot baseline. Additionally, Figure 7 provides a grid representation of accuracy variations across different pairings of real and fake exemplars.
From the results, we observe that model accuracy fluctuates depending on the selection of exemplars. The highest observed accuracy gain was +1.60%, while the largest drop was -1.00%. Despite these variations, the overall standard deviation remains moderate, indicating that while sample selection plays a role, GPT-4o exhibits relative stability in performance across different trials.
While sample selection influences performance, the variance remains limited, suggesting GPT-4o exhibits robustness to exemplar variations. In applications demanding high interpretability and reliability, exploring broader sample combinations could further optimize performance.
Figure 9. More examples with GPT-4o. Responses are summarized. Different colors of the text indicate the verdict for each query (red/green: fake/real).: